# Application of Machine Learning in Agriculture: Recent Trends and Future Research Avenues

### v1

@@ Abstract
Food production is a vital global concern and the potential for an agritech revolution through artificial intelligence (AI) remains largely unexplored. This paper presents a comprehensive review focused on the application of machine learning (ML) in agriculture, aiming to explore its transformative potential in farming practices and efficiency enhancement. To understand the extent of research activity in this field, statistical data have been gathered, revealing a substantial growth trend in recent years. This indicates that it stands out as one of the most dynamic and vibrant research domains. By introducing the concept of ML and delving into the realm of smart agriculture, including Precision Agriculture, Smart Farming, Digital Agriculture, and Agriculture 4.0, we investigate how AI can optimize crop output and minimize environmental impact. We highlight the capacity of ML to analyze and classify agricultural data, providing examples of improved productivity and profitability on farms. Furthermore, we discuss prominent ML models and their unique features that have shown promising results in agricultural applications. Through a systematic review of the literature, this paper addresses the existing literature gap on AI in agriculture and offers valuable information to newcomers and researchers. By shedding light on unexplored areas within this emerging field, our objective is to facilitate a deeper understanding of the significant contributions and potential of AI in agriculture, ultimately benefiting the research community.
@@
Agriculture , Artificial Intelligence (AI) , Machine Learning (ML) , Precision Agriculture , Smart Farming , Digital Agriculture
@@ 1 Introduction
“The future of agriculture is not about better farming, but about smarter farming.” - Unknown
@@
The dawn of human civilization can be traced back to the advent of agriculture, which has had an indelible impact on humanity. Agriculture encompasses a wide range of practices, including crop production [1] and livestock production, aquaculture, and forestry, which provide food and non-food products such as fibers, fuels, and raw materials such as rubber. It is an essential sector for feeding the growing world population, which is predicted by the Food and Agriculture Organization (FAO) to reach approximately nine billion by 2050 [2]. Agriculture also plays a critical role in the global economy, providing a foundation for sustainable growth and long-term economic development. Many countries have recognized the importance of this sector, as demonstrated by its contribution (Fig. 1) to their gross domestic product (GDP). In 2022 alone, the United States’ agricultural GDP reached \$201 billion in the fourth quarter of 2022, while China, Japan, and India also made notable contributions, reflecting the economic weight of the sector. The agricultural exports of the top five countries are depicted in Fig. 2. The increasing demand for food requires [3] continued focus on agricultural innovation and investment. To assess the global distribution of the agricultural land area, data was collected in the top $10$ countries ranked by agricultural land area in 2020 (Fig. 3). These countries not only demonstrate a significant commitment to agricultural production, but also play a crucial role in the global food supply chain.
Figure 1: GDP of top five counties [4].
Throughout history, three significant agricultural revolutions have shaped agricultural practices [5]. To enhance comprehension, a visual representation is provided in Fig. 4. The Neolithic Revolution, also known as the First Agricultural Revolution, emerged around 10,000 BCE in the Middle East’s Fertile Crescent. It witnessed the domestication of plants and animals and the rise of agriculture as a way of life. The Second Agricultural Revolution, which occurred in the $18^{th}$ and $19^{th}$ centuries in Europe and North America, introduced technological advances such as plow, seed drilling, and crop rotation, significantly increasing productivity. The Third Agricultural Revolution, known as the Green Revolution, was a global phenomenon of the mid-$20^{th}$ century. It involved the development of high-yielding crop varieties, the increase in the use of fertilizers and pesticides, the mechanization and irrigation, all of which aimed to improve agricultural productivity. Now, the fourth agricultural revolution, Agriculture $4.0$, is underway, with digital advancements and AI [6] playing a crucial role, akin to their impact on the fourth industrial revolution according to experts [7, 8].
Figure 2: Export in agriculture of top five counties [9] [10] [11].
@@
ML is an interdisciplinary field that merges computer science and statistics to perform tasks that are typically performed by humans. It relies on data analysis and classification to train machines without explicit programming. In recent years, the integration of ML with big data technologies and high-performance computing has gained popularity in agriculture, facilitating the analysis of data-intensive processes. ML has become an essential tool in agriculture and has found applications in various scientific fields such as bio-informatics [12], biochemistry [13], medicine [14], meteorology [15, 16], economic sciences [17, 18], robotics [19], aquaculture [20], food security [21], optimization [22, 23], industry [24, 25], and climatology [26].
Figure 3: Agricultural land world wise, sq. km., 2020 - Country rankings [27].
@@
Agriculture 4.0 represents a new era of farming, merging traditional methods with technologies such as IoT, AI, and automation. Through the use of sensors, drones, and AI-driven machinery, this approach increases efficiency and sustainability, enhancing yields, and reducing waste. It is a direct response to growing food needs and environmental challenges, reflecting the agricultural sector’s readiness to embrace innovation and adapt to a rapidly changing world. Precision agriculture [28], digital agriculture [29], and smart agriculture [30] are terms used to describe the use of technology and data in agricultural practices. These evolving agricultural practices hold great promise for improving efficiency, sustainability, and productivity in agriculture. Although they share some similarities, there are subtle differences in their focus and approach as follows:
  • Precision agriculture: This primarily aims to optimize the use of inputs such as water, fertilizers, and pesticides to improve crop yields.
  • Digital agriculture: This focuses on using digital technologies such as sensors, satellite imagery, and data analytics to collect and analyze data for better decision-making.
  • Smart agriculture: This involves the use of advanced technologies such as AI, robotics, and the IoT to automate and optimize agricultural processes.
Figure 4: Revolution in agriculture and key characteristics.
@@
However, the agriculture sector faces challenges that include climate change, limited resources, and the need for sustainable practices. Unpredictable weather patterns, pests, and diseases affect crop yields. Limited land, water and energy resources strain production. ML offers innovative solutions to improve productivity and address these challenges. Although some papers have explored various ML models, a comprehensive review is still needed in this emerging field. This paper aims to provide that review, focusing on the unique features and applications of recent ML models in agriculture.
The organization of the paper is as follows: Section 2 begins by outlining the scope of the studies covered in this article. Section 3 serves as an overview of ML. Section 4 shifts the focus to agriculture, explaining its various steps. Section 5 reviews recent applications of ML in different stages of farming. A quantitative analysis is presented in Section 6. Few case studies are discussed in Section 7. The transformative influence of the top $10$ AI and robotics companies in agriculture is explored in Section 8. Section 9 delves into the challenges associated with the use of AI and ML in agriculture. Looking ahead, Section 10 presents future expectations within this field. The paper comes to its conclusion in Section 11, combining the key findings and implications of the study.
@@ 2 Scope of Studies
The field of agricultural AI has seen a significant surge in research, resulting in numerous reviews and state-of-the-art research articles published over time. In Table 1, we provide a summary of some of these review articles, highlighting their focused areas and key comments. In this review article, our aim is to address the gaps observed in the existing literature to offer the best understanding of the current research landscape in agricultural AI. Our methodology consisted of a three-stage process: identifying relevant studies, screening them, and evaluating their eligibility. This approach enabled us to outline the distinct features, advantages, and shortcomings of each application, thus illuminating both the potential benefits and challenges inherent in the deployment of ML in agricultural settings. Unlike previous surveys, this review offers a comprehensive look at practical agricultural applications of top companies $10$ in AI and robotics or start-ups in agriculture and how academic research can drive innovation.
This paper aims to explore the following key areas:
  1. RO1:
    Assessing the effects and results of applying AI and ML technologies in agriculture.
  2. RO2:
    Investigating the latest developments and research in AI technologies within the agricultural industry.
  3. RO3:
    Gaps between academic research and practical industry implementation of AI in agriculture.
Table 1: Summary of Review Papers on AI in Agriculture.
Analysis Type of Review
Ref. Pesticide Mngmnt Weed Detect Diseases Detect Crop Recmd. Sys. Auto. Irrig. Control Yield Mngmnt Soil/Seed Systematic Bibliogr. Comprehen. Mixed Rev.
[31]
[32]
[33]
[34]
[35]
[36]
[37]
[31]
This paper
@@ 3. RO3:
However, our review paper does not focus on the technical intricacies of individual ML algorithms, nor does it extensively explore the broader socio-economic impacts of agricultural AI. Instead, our objective is to present a holistic and valuable analysis that offers readers a comprehensive understanding of the current state and future potential of ML in the transformation of agriculture.
3 Machine Learning: An Overview
ML involves a learning process in which algorithms acquire knowledge from training data to perform tasks. It is a field of AI focused on developing algorithms that learn from data to make predictions or decisions. By training models with data, ML enables accurate predictions or decisions with new data. Pioneered by Arthur Samuel [38] and Frank Rosenblatt in the mid-twentieth century, ML has advanced with the availability of powerful computers and extensive datasets. With the exponential growth of data in recent years, ML has become invaluable for organizations to make informed decisions, improve products and services, and optimize operations. ML finds applications in diverse industries such as finance, healthcare, transportation, and entertainment [39]. The three main types of ML are shown in Fig. 5.
Figure 5: Major classification of ML.
3.1 Supervised learning
Supervised learning is a type of ML in which a model is trained on a labeled data set, which means that the data set includes both input data and the corresponding output labels [40]. The objective of supervised learning is for the model to learn to map inputs to outputs accurately so that it can make accurate predictions or classifications when presented with new unseen data. The training process involves using various algorithms to analyze the labeled data set and develop a model that can accurately generalize and make predictions on new data. The details and key features of popular supervised learning models, including Linear Regression, Random Forests, Support Vector Machines, and Artificial Neural Networks, are presented in Table 2.
Table 2: Comparison of supervised learning Models.
SL Model Key Features Pros Cons
1 Linear Regression Simple to implement, easy to interpret Fast, works well with linearly separable data Limited to linear relationships
2 Logistic Regression Can handle binary and multi-class classification, interpretable coefficients Fast, works well with linearly separable data Limited to linear decision boundaries
3 Decision Trees Non-parametric, can handle both categorical and continuous data, easy to interpret Can capture complex non-linear relationships, robust to outliers Can overfit to the training data
4 Random Forests Ensemble of decision trees, reduces over-fitting, handles high-dimensional data Can handle complex non-linear relationships, robust to outliers Computationally expensive
5 Support Vector Machines (SVMs) Can handle high-dimensional data, works well with non-linearly separable data Can capture complex non-linear relationships Computationally expensive, sensitive to parameter settings
6 Naive Bayes Classifiers Simple and fast, can handle high-dimensional data Requires strong independence assumptions, may not work well with correlated features Limited to linear decision boundaries
7 Artificial Neural Networks (ANNs) Can learn complex non-linear relationships, can handle large amounts of data Can handle complex non-linear relationships Computationally expensive, requires careful tuning
8 K-Nearest Neighbor (KNN) Classifiers Simple and fast, non-parametric, can handle multi-class classification Can handle complex non-linear relationships, robust to outliers Computationally expensive at test time
9 Gradient Boosting Machines (GBMs) Ensemble of decision trees, handles heterogeneous data, reduces bias Can handle complex non-linear relationships Computationally expensive, sensitive to parameter settings
10 Convolutional Neural Networks (CNNs) Designed for image data, can learn hierarchical representations State-of-the-art performance on image classification tasks Computationally expensive, requires large amounts of data
3.2 Unsupervised learning
Unsupervised learning is an ML technique in which the model is trained on a data set without any predefined labels or target variable [40]. The goal is to identify patterns and relationships within the data, such as clusters or groups, and discover meaningful insights from the data. Unsupervised learning algorithms do not have a clear objective or specific output to predict, but rather aim to discover the underlying structure and relationships within the data. Examples of unsupervised learning algorithms include clustering, dimensionality reduction, and anomaly detection. The key features, pros, and cons of the most popular models in unsupervised learning are summarized in Table 3.
Table 3: Comparison of unsupervised learning models.
SL Model Key Features Pros Cons
1 Clustering Groups data points based on similarity, useful for exploratory data analysis Can identify unknown patterns, useful for outlier detection Can be sensitive to initialization, difficult to interpret results
2 Dimensionality Reduction Reduces the number of features in data, useful for visualization and data compression Can improve model performance by removing noise and irrelevant features May lose important information, difficult to interpret results
3 Association Rule Learning Identifies relationships between items in large datasets, useful for market basket analysis Can uncover hidden patterns and associations Limited to categorical data, may produce spurious associations
4 Anomaly Detection Identifies unusual or rare data points, useful for fraud detection Can identify unknown patterns and anomalies Difficult to determine appropriate threshold for anomaly detection
5 Generative Models Learns the underlying probability distribution of data, useful for data generation and synthesis Can generate new samples of data, useful for data augmentation Computationally expensive, difficult to train and interpret
3.3 Reinforcement learning
Reinforcement learning can be described as a category of ML in which an agent acquires the ability to act within an environment by taking specific actions and receiving rewards or punishments in response [40]. The goal of the agent is to learn to make the best decisions possible in order to maximize its rewards over time. The agent receives feedback from the environment in the form of rewards, which are positive or negative signals indicating whether the agent’s actions were good or bad. As the agent continues to interact with its environment, it gradually gains knowledge of which actions result in the greatest rewards and which should be avoided. This iterative learning process is commonly employed in a variety of applications, including gaming, robotics, and decision-making in intricate settings. Table 4 summarizes the key features, pros, and cons of the most popular models in reinforcement learning.
Table 4: Comparison of reinforcement learning models.
SL Model Key Features Pros Cons
1 Q-Learning Model-free, TD-based, can handle large state spaces Simple to implement, can handle complex decision-making tasks Can be slow to converge, requires careful selection of learning rate and exploration rate
2 SARSA Model-free, TD-based, can handle large state spaces Can handle dynamic and stochastic environments Can be slow to converge, requires careful selection of learning rate and exploration rate
3 Deep Q-Networks (DQN) Uses deep neural networks to approximate Q-values, can handle high-dimensional state spaces Can handle complex environments with high-dimensional state spaces Can be unstable during training, requires careful selection of hyperparameters
4 Policy Gradients Model-free, policy-based, can handle continuous action spaces Can handle complex environments with continuous action spaces Can be slow to converge, requires careful selection of learning rate and exploration rate
5 Actor-Critic Model-free, combines value-based and policy-based methods, can handle continuous action spaces Can handle complex environments with continuous action spaces Can be slow to converge, requires careful selection of learning rate and exploration rate
6 Proximal Policy Optimization (PPO) Model-free, policy-based, can handle continuous action spaces Can handle complex environments with continuous action spaces, more stable than policy gradients Can be slow to converge, requires careful selection of learning rate and exploration rate
7 Monte Carlo Tree Search (MCTS) Model-based, tree-based search, can handle large state spaces Can handle complex decision-making tasks, can make decisions based on incomplete information Computationally expensive, may not work well in continuous state or action spaces
4 Agriculture: An Overview
Agriculture is the cultivation of crops, livestock, and other natural resources for human consumption or use. It has been an integral part of human civilization since ancient times. The major branches of agriculture, including agronomy, horticulture, forestry, livestock, fisheries, agriculture engineering, and economics.
Agriculture tasks can be categorized into three main sub-areas: pre-harvesting, harvesting, and post-harvesting. Fig. 6 provides a detailed depiction of these sub-domains of agricultural tasks. During pre-harvesting tasks, the selection of suitable crops for the climate and soil conditions of the area is focused on by farmers. The land is prepared, seeds are sown, and the crops are maintained through techniques like irrigation, pesticide use, pruning, and other practices. Estimating the yield is also an important aspect of pre-harvesting activities, involving activities such as yield mapping and fruit counting to predict production and make appropriate arrangements for harvesting or post-harvesting. The process of collecting mature crops or fruits from the fields, known as harvesting, is focused on ensuring that the crops meet market quality standards. Post-harvesting activities, which include storage and processing systems, play a crucial role in maintaining the quality of the harvested produce. In each stage of farming, important factors such as soil fertility, irrigation, and pest control need to be considered.
Figure 6: Overview of the stages and steps involved in agriculture.

### v2

# Information Fusion in Smart Agriculture: Machine Learning Applications and Future Research Directions

@@ Abstract
Machine learning (ML) is a rapidly evolving technology with expanding applications across various fields. This paper presents a comprehensive survey of recent ML applications in agriculture for sustainability and efficiency. Existing reviews mainly focus on narrow subdomains or lack a fusion-driven perspectives. This study provides a combined analysis of ML applications in agriculture, structured around five key objectives: (i) Analyzing ML techniques across pre-harvesting, harvesting, and post-harvesting phases. (ii) Demonstrating how ML can be used with agricultural data and data fusion. (iii) Conducting a bibliometric and statistical analysis to reveal research trends and activity. (iv) Investigating real-world case studies of leading artificial intelligence (AI)-driven agricultural companies that use different types of multisensors and multisource data. (v) Compiling publicly available datasets to support ML model training. Going beyond existing previous reviews, this review focuses on how machine learning (ML) techniques, combined with multi-source data fusion (integrating remote sensing, IoT, and climate analytics), enhance precision agriculture by improving predictive accuracy and decision-making. Case studies and statistical insights illustrate the evolving landscape of AI driven smart farming, while future research directions also discusses challenges associated with data fusion for heterogeneous datasets. This review bridges the gap between AI research and agricultural applications, offering a roadmap for researchers, industry professionals, and policymakers to harness information fusion and ML for advancing precision agriculture.
@@
Artificial Intelligence (AI), Machine Learning (ML), Agriculture , Precision Agriculture, Smart Agriculture, Digital Agriculture , Data Fusion
[…]
[mfs]organization=Mehta Family School of Data Science and Artificial Intelligence, Indian Institute of Technology Roorkee, city=Roorkee, country=India
[…]
[ISI]organization=Statistical Quality Control and Operations Research Unit, Indian Statistical Institute, city=Hyderabad, country=India
[…]
[ie]organization=Department of Applied Mathematics and Scientific Computing, Indian Institute of Technology Roorkee, city=Roorkee, country=India
[…]
[apm]organization=Department of Industrial Engineering, University of Chile, city=Santiago, country=Chile
[…]
[math]organization=Department of Mathematics, Indian Institute of Technology Roorkee, city=Roorkee, country=India
{highlights}
Comprehensive review of ML applications in pre-harvest, harvest, and post-harvest.
Emphasizes multi-source data fusion for improved precision farming and decision-making.
Includes bibliometric and statistical analysis of ML research trends in agriculture.
Explores real-world AI applications and industry innovations in smart agriculture.
Presents key datasets supporting ML model training for agricultural advancements.
@@ 1 Introduction
Artificial intelligence (AI) is a powerful technologies with applications in various fields, such as healthcare, agriculture, cybersecurity, finance, transportation, and many more. Recently, machine learning (ML) has made significant advancements in agricultural applications, driving innovation and efficiency in the field. Agriculture encompasses a wide range of practices, including crop monitoring, weather predictions for crop yield, livestock production, aquaculture, and forestry, which provide food and non-food products such as fibers, fuels, and raw materials like rubber [1]. It is an essential sector for feeding the growing world population, which is predicted by the Food and Agriculture Organization (FAO) to reach approximately nine billion by 2050 [2]. Agriculture also plays a critical role in the global economy, providing a foundation for sustainable growth and long-term economic development. Many countries have recognized the importance of this sector, as demonstrated by its contribution to their gross domestic product (GDP). In 2022 alone, the United States’ agricultural GDP reached \$201 billion in the fourth quarter of 2022 [3], while China [4], Japan [5] and India also made notable contributions, reflecting the economic weight of the sector. The increasing demand for food requires continued focus on agricultural innovation and investment [6].
Throughout history, four major agricultural revolutions have transformed farming. The First Agricultural Revolution (Neolithic, $\sim$10,000 BCE) introduced domestication of plants and animals. The Second Agricultural Revolution (18th–19th centuries) in Europe and North America brought innovations like the plow, seed drilling, and crop rotation, boosting productivity. The Third Agricultural Revolution (Green Revolution, mid-20th century) introduced high-yield crops, fertilizers, pesticides, mechanization, and irrigation. Now, the fourth agricultural revolution; Agriculture 4.0 is underway, driven by AI, multi-sensor data fusion, and hybrid architectures, optimizing precision farming with diverse real-world datasets [166, 210, 112].
@@
ML is an interdisciplinary field that merges computer science and statistics to perform tasks that are typically performed by humans. In recent years, the integration of ML with big data technologies and high-performance computing has gained popularity in agriculture, facilitating the analysis of data-intensive processes. Also, to enhance the accuracy of ML applications and predictions, researchers and practitioners integrate data from multiple sources, a process known as information fusion [37]. ML and data fusion have become an essential tool in agriculture and have found applications in various scientific fields such as bio-informatics [104], biochemistry [165], medicine [106], meteorology [107], economic sciences [105], robotics [111], aquaculture [109], food security [108], and climatology [110].
@@
Agriculture, like other domains, is being transformed by ML. Agriculture 4.0 merges traditional practices with IoT, AI, and automation, utilizing data fusion from sensors, drones, and AI-driven machinery to enhance efficiency, sustainability, yields, and waste reduction. It is a direct response to growing food needs and environmental challenges, reflecting the readiness of the agricultural sector to embrace innovation and adapt to a rapidly changing world. Precision agriculture [113], digital agriculture [114], and smart agriculture [115] are terms used to describe the use of technology and data fusion in agricultural practices. These evolving agricultural practices hold great promise for improving efficiency, sustainability, and productivity in agriculture. The study [7] is the earliest known work in the literature to introduce AI in agriculture, where expert systems were developed as decision aids for integrated crop management, encompassing areas such as irrigation, fertilization, weed control, and pest management. These expert systems aimed to enhance decision-making in agricultural practices by using AI-based reasoning and analysis to improve efficiency and productivity.
@@
However, the agriculture sector faces challenges that include climate change, limited resources, and the need for sustainable practices. Unpredictable weather patterns, pests, and diseases affect crop yields. Limited land, water, and energy resources strain production. ML offers innovative solutions to improve productivity and address these challenges. Although some review articles (details are provided in Section 2) have explored various ML models, a comprehensive review is still needed in this emerging field. Consequently, this study aims to explore the following research objective (RO):
  1. RO-1:
    Investigating recent advancements and applications of ML in pre-harvesting, harvesting, and post-harvesting processes, highlighting early disease detection, weather prediction, seed testing, irrigation planning, and crop recommendation.
  2. RO-2:
    Demonstrating how ML can be used with agricultural data and data fusion to provide useful insights.
  3. RO-3:
    Understanding current research activity and conducting a bibliometric analysis by accumulating statistical data from the Scopus directory.
  4. RO-4:
    Examining real-world applications of the fusion of ML and agriculture, including case studies of the top 10 AI-based companies based on valuation that address the gap between academia and practical implementation.
  5. RO-5:
    Listing popular datasets for training ML models, providing resources for practitioners.
The structure of the paper is illustrated in Fig. 1, which outlines the sections and subsections.
Figure 1: The organization of this paper with sections and subsections.
@@ 2 Scope of Studies
The field of agricultural ML has seen a significant increase in research, resulting in a substantial volume of research articles. In the last two years alone, more than 5,000 articles have been published involving the keywords “agriculture” and “machine learning.” (A detailed quantitative analysis of this trend is provided in Section 5, particularly in Fig. 4.) Therefore, conducting a comprehensive systematic review on the ML and agriculture domain seems nearly impossible. This is why most available systematic reviews focus on specific niches within agriculture, such as soil classification [231], disease detection [232], or crop recommendation [234] using ML.
Although some comprehensive review articles are available on the general domain of ML and agriculture, they are not up-to-date and do not cover all aspects. A summary of some of these review articles, highlighting their focus areas and key contributions, is provided in Table 1. Therefore, by focusing on RO-1 to RO-5, this study provides a meaningful up-to-date survey of the fusion between ML and agriculture. Additionally, given the critical role of multiple data sources—such as remote sensing data (satellite, aerial, drone, LiDAR, and radar data)—this study emphasizes data fusion and its impact on agricultural advancements.
Our methodology consisted of a three-stage process: identifying relevant studies, screening them, and evaluating their eligibility. This approach allowed us to outline the distinct features, advantages, and shortcomings of each application, thus highlighting both the potential benefits and the challenges inherent in the deployment of ML in agricultural settings. Relevant data has been collected from the Scopus database.
Table 1: Summary of Review Papers on Machine Learning in Agriculture.
Analysis Type of Review Data Fusion
Ref. Pesticide
Management
Weed
Detection
Disease
Detection
Crop Recommen-
dation System
Automatic
Irrigation Control
Yield
Management
Soil/Seed
Analysis
System-
atic
Bibliog-
raphy
Compre-
hensive
Mixed Data Fusion Covered
[194] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
[191] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
[190] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
[189] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
[192] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
[197] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
[188] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
[8] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
This paper $\times$
@@
However, this comprehensive review does not focus on the technical intricacies of individual ML algorithms, nor does it extensively explore the broader socio-economic impacts of agricultural AI. Instead, our objective is to present a holistic and valuable analysis that offers readers a comprehensive understanding of the current state and future potential of ML in the transformation of agriculture.